OpenStudy (anonymous):

A circle has a diameter of ten yards. What is its area? Use three point one four for pi thirty-one point four square yards thirty-nine point two five square yards sixty-two point eight square yards seventy-eight point five square yards

OpenStudy (anonymous):

\[A = \pi r^2 \] r = d/2 r = 5 \[A = \pi 5^2\] \[A = 25 \pi\] A = 78.5 Answer is D.